How far does your car, moving at 55 mil/h travel forward during the one second of time that you take to look at an accident on the side of a road?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Shahid
 
It means that car travel 55 mikes in 1 hour.
Hence
In 60 minutes car travels = 55 miles
1 min = 55/60 miles
60 sec = 55/60 miles
 
Hence in 1 sec  = 55/3600 miles
 = 0.01527 miles
 
Regards
Arun (askIITians forum expert)